# DF40: Toward Next-Generation Deepfake Detection

### v1

@@ 4.1 Experimental Setup
All pre-processing and training codebases in this work adhere to DeepfakeBench [94] to align the standardized settings. By following these standardized procedures, we aim to provide reliable and reproducible results for evaluations. The details of dataset configuration, algorithm implementation, and full training can be seen in the supplementary. Regarding evaluation, we compute the frame-level Area Under the Curve (AUC) as our primary evaluation metric. For detection selection, we choose the classical baseline Xception [16] with four SoTA detectors (SRM [52], SPSL [48], RECCE [8], RFM [80]) that use Xception as the backbone and implement a new algorithm for improved detection. In this manner, we aim to revisit the extra improvement over the baseline under our setting. We also include one SoTA blending-based method SBI [70]. Note that the original SBI paper only uses "BlendFake" (BI) for training, so here we do not train it with DF40 in evaluations.
Figure 4: Full results of SBI [70] on the proposed DF40 dataset for both the FF (within) and CDF domain (cross). Discussion and analysis for the figure can be seen in Sec. 5.1. We evaluate other detectors, including Xception [67] and SPSL [48]) on our DF40 (CDF) in Fig. 9.
4.2 Evaluations
@@
In this work, we conduct main evaluations using four types of protocols: Cross-forgery evaluation (Protocol-1), Cross-domain evaluation (Protocol-2), Toward unknown forgery and domain evaluation (Protocol-3), and One-Verse-All (OvA) evaluation (Protocol-4). Our findings are highlighted in blue. We show the observations and use t-SNE, logit distributions, and the causal graph for analysis.
Protocol-1: Cross-forgery evaluation.
This evaluation aims to evaluate the model’s performance within the same data domain but across different forgery methods. The results are presented in Tab. 4, and the observations are as follows: Finding-1: Asymmetric performance drop among different forgery types. Notably, the performance drop between FS and FR is moderate but more significant between these two and EFS. Specifically, model training on FS yields higher results (around 0.8) on EFS, while training on EFS only achieves about 0.6 on FS. Finding-2: Existing well-designed (SoTA) detectors may not have obvious advantages over the baseline. In our experiments, SRM, SPSL, RECCE, and RFM, all state-of-the-art deepfake detectors using Xception as their backbone, achieve results similar to the simple baseline Xception. This implies that these SoTA models might only learn sub-optimal forgery features and lack clear advantages over the baseline, although these SoTAs show higher results on previous datasets, $e.g.$, DFDC. Finding-3: CLIP excels in deepfake detection than other baselines. Outperforming other SoTA detectors and the Xception baseline across all scenarios, CLIP demonstrates the power of pre-training in deepfake detection. This raises the question: can we perform pre-training in the specific deepfake domain instead of the general CV domain? We will explore pre-training with CLIP in Sec. 5.2. Finding-4: Blending is not all you need to detect deepfakes, even for face-swapping. As observed, the SoTA blend-based detector SBI only achieves an average AUC result of 0.734, although it has demonstrated significant generalization performance on previous deepfake datasets. Previous research [45; 70; 12] has shown that models trained on blending data can achieve significantly higher results than those trained on deepfake data, suggesting that using blending for training is a necessary condition to achieve SoTA results. However, our results indicate that blending is not all you need. We conduct further analysis by testing SBI on individual forgery types of our DF40 to determine which type of forgery SBI can successfully simulate (see Sec. 5.1).
Protocol-2: Cross-domain evaluation.
@@
Cross-domain evaluation assesses the model’s performance across different data domains but with the same forgery methods. The results are presented in Tab. 4, and the observations are as follows: Finding-5: Relatively minor performance drop when solely crossing data domains or forgery methods. As we can see, the performance drop between FS and EFS is relatively small (0.9+) compared to FR (0.8+). Also, we found that the discriminative forgery artifacts of EFS do not have solid casual with data domain a very minor performance drop (only about 1%). Finding-6: Forgery methods and data domains together contribute to discriminative forgery artifacts. A significant performance drop is observed when both forgeries and domains change. For example, Xception training on FS (FF) achieves only 0.657 and 0.642 on FR (CDF) and EFS (CDF), respectively. These results are significantly lower (nearly a 20% drop) compared to results on FS (CDF) and FR (FF), which involve crossing only one of the domains or forgeries. This suggests that both factors together contribute to discriminative forgery artifacts for distinguishing real and fake.
Figure 5: Causal graph for analyzing the influence of domain (D) and forgery (F).
@@
We conduct further casual analysis below for better clarification. To intuitively convey our conclusions, we draw a causal graph to illustrate the causal relationship of how domain (D) and forgery method (F) influence the model’s generalization (R). The causal graph is presented in Fig. 5. In this graph, D represents the data domain, and F represents the forgery method. We hypothesize that these two factors influence three intermediate variables, $X_{1}$, $X_{2}$, and $X_{3}$, which in turn affect the result $R$. The relationships are described as follows: $X_{1}$ represents the effect of the domain (D) on the result (R) independently; $X_{3}$ represents the effect of the forgery method (F) on the result (R) independently; and $X_{2}$ represents the combined effect of both the domain (D) and the forgery method (F) on the result (R). Mathematically, the model’s generalization result $R$ can be expressed as a function of these intermediate variables: $R=f(X_{1},X_{2},X_{3})$, where $X_{1}=g(D)$ represents the domain-specific influence, $X_{3}=h(F)$ represents the forgery method-specific influence, and $X_{2}=k(D,F)$ represents the combined influence of both domain and forgery method. Each of these variables $X_{1}$, $X_{2}$, and $X_{3}$ captures different aspects of how D and F interact to impact the overall performance $R$. The functions $g$, $h$, and $k$ are mappings that quantify the respective influences: $X_{1}=g(D)$, $X_{3}=h(F)$, and $X_{2}=k(D,F)$. Finally, the result $R$ is determined by integrating these influences: $R=f(g(D),k(D,F),h(F))$. This formalization helps us understand and analyze the separate and combined effects of the data domain and forgery method on the model’s generalization capability.
The logits and confidence analysis (shown in Fig. 6) further supports this point, showing the increased overlap between real and fake logits as domains and methods change incrementally. Also, crossing domains can result in "wrong confidence" for real predictions and vice versa. More specifically, we can observe from Fig. 6, as follows: (1) For Xception: (i) Same Domain and Method: Real and fake class distributions have small variance and high mean, with minimal ambiguity in confidence and less overlap between the real and fake logits. (ii) Cross Method, Same Domain: Fake class distribution has increased variance and lower mean, while real class remains relatively unchanged. (iii) Cross Domain, Same Method: Real class distribution has increased variance and lower mean, while fake class remains relatively unchanged, causing a slight performance decrease. (iv) Cross Method and Domain: Both real and fake class distributions have significantly increased variance and lower means, decreasing overall performance. Furthermore, in this case, the overlap between the real and fake distributions becomes obviously larger than the case(i), where both the domain and method are the same. (2) For SBI: It shows a different behavior than Xception. The real class distribution consistently has a small variance and high mean, while the fake class distribution has a larger variance and lower mean. This might suggest that SBI can learn a more robust distribution for real classes than the baseline (wo using blending), generalizing well in the FF domain using the FS method (0.81 AUC). However, when the real class distribution changes ($e.g.$, in the CDF domain), performance drops significantly (to 0.679 AUC). We will give a further discussion of SBI in Sec. 5.1.
Protocol-3: Unknown forgery and domain evaluation.
This protocol tests model performance across different data domains and forgery methods, incorporating SoTA and popular software like DeepFaceLab [18], MidJourney-6 [53], and HeyGen [28] to simulate real-world scenarios. From the whole results in Tab. 5, we can see that: (i) Forgery artifacts can vary significantly across different forgeries (especially for FR and EFS), and artifacts of EFS can be similar to that of FE (see the results of StarGAN2 and CollabDiff); (ii) Models trained on FR display limited generalization ability, even falling below chance levels (0.5 AUC). Notably, Xception achieves only a 0.006 AUC on StyleCLIP. We provide further analysis of this phenomenon with both logit distribution and t-SNE visualization in Sec. 5.9. (ii) HeyGen and MidJourney-6 seem to pose significant challenges for detecting unknown forgeries, as most detectors only achieve unexpected results (AUC below 0.5). A possible reason for this is that both testing datasets exhibit highly realistic forgery quality. Consequently, it is essential to develop more robust detection methods in the future that can effectively identify forgeries with not only high-quality fakes but also high-resolution images.
Figure 6: Logits and confidence distribution analysis. We visualize the logits distribution (before SoftMax) and confidence distribution (after SoftMax) for both real and fake classes. The overlap in the logits distribution represents the portion of "wrong classifications." We analyze two models: Xception [67] and SBI [70]. More discussion and analysis can be referred to Sec. 5.
Figure 7: t-SNE Visualizations for three models: CLIP-large, CLIP-base, and Xception (from left to right). These models are trained on FS (FF). We conduct visualizations on FS (CDF), HeyGen [28], MidJourney [53], and WhichisReal [53]. More discussion and analysis can be referred to Sec. 5.
Protocol-4: One-Verse-All (OvA) evaluation.
@@
The OvA evaluation trains a model on a specific forgery and tests it on all other forgeries. The results, presented in Fig. 3, reflect 31$\times$31 evaluations for the heatmap (excluding 9 unknown domain forgeries). Finding-7: FR forgeries may share transferable patterns for detecting other FR instances. Most FR forgery methods demonstrate high AUC within the heatmap, even across data domains. However, Wav2Lip, an audio-driven reenactment forgery that only modifies the mouth region, is an exception. Its artifacts are more localized compared to other FRs that generate all content, making it different.
@@
Table 7: Similar to Tab. 6, we evaluate on GenImage [102] (non-face domain), pure AIGCs.
@@
5 Analysis
In this subsection, we will dive deeper to provide insightful analysis based on previous observations and offer several new findings to the detection field to inspire future research.
5.1 Is Blending All You Need for Deepfake Detection?
Blending-based detectors have been considered highly effective for generalizable deepfake detection. Many SoTA detectors [70; 45; 11] use blending-generated pseudo-fake data and exclude actual deepfake data for training. However, our evaluations empirically show that blending might not all you need, as models trained with blending generally underperform compared to those trained on actual deepfake data (Tab. 4, Tab. 4, Tab. 5). This is likely because blending simulates fake artifacts similar to the "DF-family" of Face Swapping (FS) and may struggle to generalize forgeries that do not involve blending, such as the "FS-family" of FS and AI-generated content (EFS). Also, we evaluate SBI on 31 known forgeries across FF++ and CDF domains (see Fig. 4) and find that SBI performs well on some FS forgeries but also achieves near-perfect results on non-blending methods, $e.g.$, SD-2.1. However, it performs poorly on RDDM (FF), far below chance levels. Thus, an interesting and meaningful question (Question-1) arises: Which types of forgeries can be addressed by blending? And how to effectively combine blending fakes with actual deepfakes for training? Moreover, we also found that: Finding-8: SBI could be viewed as an "anomaly detection" model. SBI can be considered an "anomaly detection" model, which might utilize the real features for classifying fakes. Results in Fig. 6 show that when forgery samples appear significantly different from real samples, SBI classifies them as anomalies (forgery samples), likely because the pseudo-fake samples generated by SBI closely resemble their original real counterparts.
5.2 Desirable Baselines/Backbones in Deepfake Detection for Pre-training.
Xception is often used as a baseline for deepfake detection, but is it the most suitable backbone? In this work, we re-evaluate Xception and compare it with the "stand-out" model CLIP (see Finding-3). We also include the large version of CLIP (CLIP-large), attempting to assess the "scaling law." As seen in Tab. 6, CLIP-large significantly outperforms CLIP-base (+8.4%) and Xception (+20.54%) on average. We also evaluate non-face domain data to determine if models trained on face-domain data can transfer to non-face-domain detection. In Tab. 7, surprisingly, CLIP-large achieves a 0.746 AUC on previously unseen non-face AIGCs, while Xception only reaches a 0.535 AUC, near chance levels. This leads to Finding-9: CLIP-large is the most powerful baseline, significantly outperforming Xception and Finding-10: CLIP-large can even generalize to some non-face AIGCs when trained only on face data. However, why can CLIP-large outperform the base by a notable margin? We analyze this point in the following paragraph from the view of real face distribution.
5.3 Analysis of the Real Face Distribution.
The results in Tab. 6 raise the question: why does CLIP-large significantly outperform Xception and its base version? We investigate further by analyzing the learned features of the three models, as visualized in Fig. 7. (1) We see that both versions of CLIP learn more informative features that can gather some real samples into several different "groups" (see HeyGen and MidJourney), while the real features of Xception mix all of them together. Leveraging large-scale pre-training, CLIP can capture more informative facial features ($e.g.$, ID) about real faces and thus "know" that these features are unrelated to deepfakes, while Xception could be overfitted to ID, as evidenced by [23]. (2) Between the large and base versions, CLIP-large’s real samples are closer to each other (see WhichisReal), suggesting it learns more comprehensive features of real faces. We also see that Finding-11: CLIP-large can implicitly "separate and align" different forgeries and data domains ("orthogonal" in WhichisReal), with no need for explicit constraints as [75]. This point is further discussed and confirmed through t-SNE and logit distribution analysis in Sec. 5.4 below.
5.4 Comparative Analysis of CLIP-base and CLIP-large for Deepfake Detection
t-SNE visualization:
Logits Distribution Analysis:
@@ 5.4 Comparative Analysis of CLIP-base and CLIP-large for Deepfake Detection
Figure 8: Frequency analysis on selected datasets of DF40. We compute the average frequency spectra of high-pass filtered images, following [82]. Although the deepfake types differ, they can show similar patterns/artifacts in the average spectra. For instance, methods from FS ($i.e.$, SimSwap [13], BlendFace [71], InSwap) show similar "checkboard" patterns that can also be observed in EFS ($i.e.$, SD1.5 [64], VQGAN [24]) and FE ($i.e.$, e4e [78]).
Figure 9: Results of evaluating three (previous) SoTA models (trained on FF++), $i.e.$, Xception [67], SPSL [48], and SBI [70]), on our DF40 (CDF). We show that, in most cases, these SoTAs fail to generalize well in our proposed forgeries, highlighting the importance of developing a comprehensive and diverse dataset.
  Model Training Set Testing Set
ADM BigGAN GLide MidJourney SD-v4 SD-v5 Vqdm Wukong Avg.
FR (FF) 0.695 0.595 0.230 0.505 0.437 0.417 0.549 0.519 0.493
EFS (FF) 0.912 0.992 0.906 0.591 0.729 0.72 0.956 0.713 0.815
DF40 (FF) 0.911 0.967 0.736 0.571 0.630 0.614 0.882 0.660 0.746
5.5 Generalization to Non-face Domain AIGCs
Table 9: Results of evaluating three models ($i.e.$, I3D, CLIP-base, and CLIP-large) trained on different variants of the proposed DF40 on other previous deepfake datasets. Note that only FF-F2F [77] and FF-NT [76] are FR methods, while other testing data all belong to FS.
  Model Training Set Testing Set
FR (FF) 0.684 0.917 0.846 0.374 0.895 0.876 0.807
FS (FF) + FR (FF) 0.764 0.970 0.926 0.994 0.928 0.953 0.929
CLIP-base FS (FF) 0.823 0.965 0.678 0.985 0.585 0.932 0.945
FR (FF) 0.649 0.669 0.572 0.457 0.682 0.625 0.947
EFS (FF) 0.553 0.618 0.604 0.634 0.539 0.653 0.590
DF40 (FF) 0.799 0.932 0.698 0.973 0.683 0.889 0.937
CLIP-large FS (FF) 0.905 0.989 0.777 0.991 0.608 0.963 0.965
FR (FF) 0.698 0.756 0.614 0.487 0.738 0.626 0.983
EFS (FF) 0.616 0.733 0.634 0.730 0.571 0.772 0.739
DF40 (FF) 0.891 0.978 0.753 0.973 0.717 0.927 0.986
5.6 Artifacts of Deepfake Forgeries in Frequency
Brief Introduction:
Mean Analysis:
Finding-12: We observe that many GAN-generated images, such as those from StyleGAN, often exhibit a checkerboard or grid-like pattern of bright spots on the mean heatmap This phenomenon occurs because of the upsampling methods used in the GAN architecture, which can introduce such artifacts during the image generation process. These patterns are typically a result of the transposed convolution operations, which can cause uneven overlaps and thus create these visible grid artifacts. Moreover, it appears that different forgery methods can share similar patterns. For instance, face-swapping techniques like SimSwap also exhibit these GAN-like checkerboard artifacts. This similarity indicates that these methods generate the entire image using AI without a blending process. This contrasts with other Deepfake methods, such as those used in Celeb-DF, which involve blending and thus do not produce the same checkerboard artifacts.
Std Analysis:
5.7 Generalization to External/Previous Deepfake Datasets
We also evaluate the model trained on DF40 to detect previous deepfake datasets. Results in Tab. 9 show that (1) For the I3D detection method, training on FS (FF) and FR (FF) combined (FS (FF) + FR (FF)) results in better performance across most testing sets compared to training on FS (FF) or FR (FF) individually. This suggests that combining face-swapping and face-reenactment datasets for training improves the generalization capabilities of the I3D method. (2) Among the CLIP-base detection methods, training on DF40 (FF) leads to better performance across most testing sets compared to training on FS (FF), FR (FF), or EFS (FF) individually. This indicates that incorporating diverse deepfake techniques in the training set improves the generalization capabilities of the CLIP-base method. (3) The same trend is observed for the CLIP-large detection methods, where training on DF40 (FF) results in better performance across most testing sets compared to training on FS (FF), FR (FF), or EFS (FF) individually. This further supports the importance of training on diverse deepfake techniques to enhance the generalization capabilities of the detection methods. (4) The CLIP-large detection method generally outperforms the CLIP-base method across most testing sets when trained on the same training sets. This suggests that the larger model size of CLIP-large contributes to better deepfake detection performance (similar to Finding-9).
5.8 Detection Results of Previous Detectors on the Proposed DF40 dataset
To evaluate the generalization capabilities of various deepfake detection models, we tested three pre-trained (on FF++) models: Xception (baseline), SPSL (frequency-based model), and SBI (blending-based model) on our DF40 (CDF) dataset, which includes 31 known deepfake methods. These models were initially trained on the FF++ dataset. The results in Fig. 9 indicate that all three models struggle to generalize effectively to the DF40 dataset, with an average AUC of only around 0.6. This performance is significantly below the ideal detection threshold, underscoring the limitations of current models in handling a diverse range of deepfake techniques. This finding highlights the critical need for developing larger and more comprehensive datasets to train robust deepfake detection systems. By expanding the diversity and volume of the training data, we can improve the models’ ability to detect and generalize across various deepfake methods, enhancing their real-world applicability and effectiveness.
Figure 10: t-SNE visualization and logits analysis for analyzing the model, which is trained on FR (FF) and tested on StyleCLIP, achieving only about 0.05 AUC in Tab. 5 (Protocol-3). See the text in Sec. 5.9 for our analysis and more details.
Figure 11: Logits and confidence analysis for comparing CLIP-base and CLIP-large models.
5.9 Analysis of "Anomalous Values" in Experiments
In our previous experimental results summarized in Tab. 5 (protocol-3), we observed an anomalous outcome where the Xception model, trained on FF (FR), achieved an AUC of less than 0.2 on the StyleCLIP dataset. Given that an AUC of 0.5 indicates performance by chance, this result is particularly perplexing. To investigate further, we conducted a t-SNE analysis on the feature distributions of different classes, as well as examined the logits and confidence distributions for different prediction results.
The t-SNE analysis 10 revealed that the features learned during training for real and fake images were inversely mapped during testing. Specifically, the features for the training real images overlapped with the testing fake images, indicating that the model might not have effectively learned robust real features. This inversion suggests that the model’s naive discriminative learning approach failed to adequately differentiate between real and fake images, leading to a complete misclassification in the new testing set.
The logits distribution 10 further supports this observation. The confidence curves for real and fake predictions were flipped, reinforcing the notion that the model was predicting the opposite of what it should. This inversion implies that the Xception model, in this context, was not only failing to generalize but was actively misclassifying due to its inability to learn stable and transferable features.
This analysis highlights a critical limitation of simple discriminative learning approaches in deepfake detection. It underscores the need for more sophisticated training strategies that can better capture the intrinsic characteristics of real and fake images, ensuring more robust performance across varied datasets. This could involve incorporating more diverse training data, using advanced feature extraction techniques, or employing generative adversarial networks to improve feature robustness and generalization.
6 Open Questions and Potential Topics for Future Research
Here, we will give further discussion on interesting research questions to inspire future research toward next-generation deepfake detection.
(1) Question-1 in Sec. 5.1 highlights the role of blending data in training deepfake detectors. Instead of using "BlendFake" only for training, how can we effectively mine the true potential and position of BlendFake? For instance, how can we combine Deepfake and BlendFake together for improved training (creating "1+1>2" results)? And what kinds of forgeries can BlendFake address?
(2) Also, as the diversity of deepfake data increasingly goes up, Question-2 raise: how to design an incremental learning framework to effectively learn (many) different forgeries incrementally? In the implementation of this incremental learning framework, how can we deal with the real data?

### v2

@@ 4.1 Experimental Setup
All pre-processing and training codebases in this work adhere to DeepfakeBench [90] to align the standardized settings. The details of dataset configuration, algorithm implementation, and full training can be seen in the Appendix. For detection selection, we choose the classical baseline Xception [12] with four SoTA detectors (SRM [48], SPSL [44], RECCE [6], RFM [75]) that use Xception as the backbone. In this manner, we aim to revisit the extra improvement over the baseline under our setting. We also consider the blending-based detector that uses the blending image (BI) for training, which creates pseudo-fakes by image blending, without using the actual deepfake data. We use SBI [65] to implement the blending-based detector. Note that the original SBI paper only uses BI for training, so we do not re-train it using our DF40 in evaluations.
4.2 Evaluations, Findings, and Analysis
@@
In this work, we conduct evaluations under four standard protocols: Cross-forgery evaluation (Protocol-1), Cross-domain evaluation (Protocol-2), Toward unknown forgery or domain evaluation (Protocol-3), and One-Verse-All (OvA) evaluation (Protocol-4). From these evaluations, we list 8 critical findings via empirical observations or in-depth analysis, as follows.
Finding-1: Asymmetric performance drop among different forgery types (fake regions matter). From Tab. 4, we observe that the performance drop between FS and FR is moderate but more significant between these two and EFS. Specifically, model training on FS yields higher results (around 0.8) on EFS, while training on EFS only achieves about 0.6 on FS. To understand the underlying reason for this finding, we provide the following detailed explanations: One possible explanation is that FS can exhibit both localized and global forgeries, while EFS is restricted to global forgeries, making EFS less diverse in scope compared to FS. As a result, training a model on EFS might not be sufficient to capture the localized fake artifacts present in FS. For example, some localized FS (such as DeepFakes of FF++ [62]) only contain fake artifacts within the facial region, but a model trained on EFS might be biased to "expect" artifacts across the entire image, including the background, resulting in a lower result on FS. To verify our claim, we consider using t-SNE for visualizing the latent distribution of ‘real’, ‘whole-fake’, and ‘face-fake’. The t-SNE (Fig. 4) results show that a well-trained EFS detection model can easily distinguish the whole-fake from real images. However, the localized version, including face-fake and mouth-fake, appears much more similar to real images, making it more challenging to detect compared to the whole-fake. This experiment qualitatively verifies the significance of the fake region in detection.
Figure 4: t-SNE visualization for real, whole-fake, face-fake, and mouth-fake images. The results show that a well-trained EFS detector (Xception) can effectively distinguish between whole-fake and real images, but struggles to identify fakes with only face or mouth manipulation. This observation highlights the significant influence of the manipulated region on detection performance.
Finding-2: Existing well-designed (SoTA) detectors may not have obvious advantages over the baseline. In our experiments, SRM, SPSL, RECCE, and RFM, all state-of-the-art deepfake detectors using Xception as their backbone, achieve results similar to the simple baseline Xception. This implies that these SoTA models might only learn sub-optimal forgery features and lack clear advantages over the baseline, although these SoTAs show higher results on previous datasets, $e.g.$, DFDC. In other words, although these well-designed SoTA detectors achieve higher scores than the Xception (baseline) when testing on CDF [42], they cannot maintain a consistent advantage over the baseline when testing on other forgeries. This highlights the remaining generalization issue.
Finding-3: CLIP excels in deepfake detection than other baselines (such as Xception). Outperforming other SoTA detectors and the Xception baseline across all scenarios, CLIP (both base and large versions) demonstrates the power of pre-training in deepfake detection. This highlights the benefits of pre-training before fine-tuning on the deepfake detection task. However, why can CLIP outperform the Xception by a notable margin? We analyze this point in the following content from the view of real face distribution. We analyze the learned features of the three models (CLIP-base, CLIP-large, Xception), as visualized in Fig. 7. (1) We see that both versions of CLIP learn more informative features that can gather some real samples into several different "groups" (see HeyGen and MidJourney), while the real features of Xception mix all of them together. Leveraging large-scale pre-training, CLIP can capture more informative facial features ($e.g.$, ID) about real faces and thus "know" that these features are unrelated to deepfakes, while Xception could be overfitted to ID, as evidenced by [19]. (2) Between the large and base versions, CLIP-large’s real samples are closer to each other (see WhichisReal), suggesting it learns more comprehensive features of real faces. We also see that CLIP-large can implicitly "separate and align" different forgeries and data domains ("orthogonal" in WhichisReal), with no need for explicit constraints as [70].
@@
Finding-4: Forgery methods and data domains together contribute to discriminative forgery artifacts. A significant performance drop is observed when both forgeries and domains change. For example, Xception training on FS (FF) achieves only 0.657 and 0.642 on FR (CDF) and EFS (CDF), respectively. These results are significantly lower (nearly a 20% drop) compared to results on FS (CDF) and FR (FF), which involve crossing only one of the domains or forgeries. This suggests that both factors together contribute to discriminative forgery artifacts for distinguishing real and fake.
Figure 5: Causal graph.
@@
To intuitively convey our conclusions, we draw a causal graph to illustrate the causal relationship of how domain (D) and forgery method (F) influence the model’s generalization (R). The causal graph is presented in Fig. 5. Mathematically, the model’s generalization result $R$ can be expressed as a function of these intermediate variables: $R=f(X_{1},X_{2},X_{3})$, where $X_{1}=g(D)$ represents the domain-specific influence, $X_{3}=h(F)$ represents the forgery method-specific influence, and $X_{2}=k(D,F)$ represents the combined influence of both domain and forgery method. Each of these variables $X_{1}$, $X_{2}$, and $X_{3}$ captures different aspects of how D and F interact to impact the overall performance $R$. The functions $g$, $h$, and $k$ are mappings that quantify the respective influences: $X_{1}=g(D)$, $X_{3}=h(F)$, and $X_{2}=k(D,F)$. Finally, the result $R$ is determined by integrating these influences: $R=f(g(D),k(D,F),h(F))$. Fig. 6(a) supports this point, showing increased overlap between real and fake logits as domains and methods change incrementally. Also, crossing domains can result in "wrong confidence" for real predictions and vice versa.
Figure 6: Logits and confidence analysis for Xception and SBI models.
Figure 7: t-SNE Visualizations for three models: Xception, CLIP-base, and CLIP-large.
@@
Finding-5: FR forgeries may share transferable patterns for detecting other FR instances. Most FR forgery methods demonstrate high AUC within the heatmap, even across data domains. However, Wav2Lip, an audio-driven reenactment forgery that only modifies the mouth region, is an exception. Its artifacts are more localized compared to other FRs that generate all content, making it different. Note that Finding-5 does not mean that collecting (many) FR methods is without value. Specifically, even though these FR methods may share some transferable patterns to detect other FR methods, the fake similarity between FR and other types (like EFS) can also vary greatly depending on the specific FR method used (see Fig. 3).
Finding-6: SBI could be viewed as an "anomaly detection" model. SBI [65] (a blending-based detection) could be considered an "anomaly detection" model, which might utilize the real features for classifying fakes. Results in Fig. 6(a) show that when forgery samples appear significantly different from real samples, SBI classifies them as anomalies (forgery samples), likely because the pseudo-fake samples generated by SBI closely resemble their original real counterparts. In other words, SBI creates more realistic fake samples that are more similar to the real, encouraging the detection model to learn a more robust real representation than the baseline.
Finding-7: CLIP-large shows the potential to generalize to some non-face deepfakes when trained only on face data. In additional to evaluate on the face deepfakes, we also evaluate non-face-domain deepfakes (natural image synthesis) using the widely-used GenImag edataset to determine if models trained on face-domain data can transfer to non-face-domain detection. In Tab. 19, surprisingly, CLIP-large achieves a 0.746 AUC on previously unseen non-face deepfakes, while Xception only reaches a 0.535 AUC, near chance levels. This highlights that CLIP-large might learn some transferable (EFS) forgery features that are not related to the face content.
@@
Table 7: Similar to Tab. 6, we evaluate on GenImage [95] (non-face domain deepfakes).
@@
Figure 8: Evaluation of existing detectors on previous datasets and our DF40.
4.3 Discussion & Further Analysis
Why A More Diverse Deepfake Dataset Is Crucial for Current Deepfake Detector?
In our work, we propose the DF40 dataset, with highly diverse deepfake approaches. Here, we aim to qualitatively verify why such a highly diverse dataset is extremely crucial for current deepfake detectors. The results in Fig. 8 show that CFDA [43] (latest SoTA) achieves significant advantages over other methods on existing datasets, but it performs comparably or even worse than other detectors. This emphasizes the necessity of constructing a dataset that encompasses a greater variety of deepfake methods and types. Without such a highly diverse dataset for evaluation, we would only be seeing “the tip of the iceberg." In other words, the current SoTA may not be the definitive “winner."
Are Super-Resolution Images Fake?
Here, we explore whether a super-resolution (SR) image will be considered fake or real by the detection model. This question is important for two reasons: (1) many deepfake software programs involve an SR operation to enhance the resolution of generated faces (e.g., FaceFusion [^3] https://github.com/facefusion/facefusion/blob/master/facefusion/processors/modules/face_enhancer.py), and (2) Most existing SR methods [81] are based on deep generative models such as GAN. Therefore, the output of SR methods can also contain generative artifacts that might be similar to those in deepfake images. We use GFPGAN to perform SR on the self-reconstruction images (SRI), where we use SimSwap [9] to perform face-swap to the same ID. Results in Tab. 8 show that SR has a significant impact on fake images, particularly for EFS and FE methods. As demonstrated in the table, models trained on FS and EFS exhibit more than a 20% point improvement for the SRI after applying super-resolution. This suggests that SR operations may introduce noticeable generative artifacts that can be detected by models trained on EFS and FE.
Exploration of Frequency Artifacts in Different Types of Deepfakes.
@@ 4.3 Discussion & Further Analysis
Here, we explore whether there exist common fake patterns in the frequency domain. following [77], we compute the average frequency spectra of high-pass filtered images. Results in Fig. 9 show that although the deepfake types differ, they can show similar patterns/artifacts in the average spectra. For instance, methods from FS ($i.e.$, SimSwap [9], BlendFace [66], InSwap) show similar "checkboard" patterns that can also be observed in EFS ($i.e.$, SD1.5 [59], VQGAN [20]) and FE ($i.e.$, e4e [73]). This observation highlights the need to further explore whether and why forgeries with different types can exhibit similar patterns in frequency.
A Model Bias Toward the Resolution Gap Between Real and Fake.
In our previous experiments (see Tab. 5), we observe anti-intuitive phenomena: models training on FR achieve very low AUC (much lower than 0.5) on StyleCLIP. After investigating the underlying reason, we found that the unaligned resolution between the real and fake classes introduces a model bias, that is, higher resolution, higher probability to be real. Specifically, we observe that the training real data (FF++) obviously contains more high-frequency components (higher resolution) compared to the training fake data (FR). This noticeable resolution gap between real and fake data could be inevitably captured by the model [56], leading the model to potentially adopt a trivial solution: "Low resolution implies fake." In contrast, during testing, the fake data (StyleCLIP) contains significantly more high-frequency components (higher resolution) than the testing real data used for testing. We provide a frequency visualization as evidence to validate this claim (see our Appendix). This bias leads the model to make an inconsistent decision, resulting in the anomaly in AUC.
Table 8: Ablation study regarding the impact of super-resolution to the fake images. We use self-reconstruction images (SRI) as the fake and apply GFPGAN [81] to perform the super-resolution.
  Test / Train FS FR EFS FE
FSGAN BlendFace LIA Wav2Lip DiT DDIM e4e
SRI 0.772 0.835 0.746 0.564 0.687 0.713 0.543
SRI + Super-Resolution 0.983 0.825 0.988 0.833 0.997 0.946 0.978
Figure 9: Frequency analysis on selected fake data within the proposed DF40.
4.4 Open Questions and Potential Topics for Future Research
(1) Regrading Blending: Many existing detectors have utilized blending data to enhance the model’s generalization. However, we have found that even the latest blending detector [43] is still limited in generalizing to all types of deepfakes (see Fig. 8). This indicates that blending might not be "all you need." But the question remains: What is the role of blending data in training deepfake detectors? And what kinds of deepfakes can be addressed by blending? (Question-1).
(2) Regrading Forgery Diversity: As the diversity of deepfakes increasingly goes up, Question-2 raise: how to design a new framework to learn (many) different forgeries effectively and jointly, without overfitting to limited specific fakes?